OpenStudy (mathmate):

Soil technology @ASH90 see comments below.

OpenStudy (mathmate):

Your original question: hi, no it's not a lab it's kind of like how we did the others but this one seems to be a bit different because on the last one we did not have to find a moist unit weight but with this one i do. So could you please check what i have as my answers for these to see if i got them correct? For a given soil sample, the moist soil sample has a volume of 0.25 ft 3 and weighed 30.75 lbs moisture content (w)= 9.8% and apparent specific gravity (Gs)= 2.66 determine a moist unit weight. V= Volume (W)= Moisture content (Gs)=Specific gravity W=Weighed ------------------------------------------------------- (a) moist unit weight=123 lb/ft^3 (b) dry unit weight= 112lb/ft^3 (c) void ratio=0.48 (d) porosity=.324

OpenStudy (mathmate):

My reply (your answers are all correct) (a) moist unit weight: correct as you had it. Moist unit weight = \(Ww/vol=30.75/0.25 lb/ft63\) (b) dry unit weight \(\gamma_d\) : correct as you had it. Moisture content W=\(\dfrac{\gamma_w-\gamma;_d}{\gamma_d}\times 100%\) so \(\gamma_d = \dfrac{\gamma_w}{1+W}\) \(= 123/(1+0.098)\) \(= 112.0~~lb/ft^3\) (c) voids ratio = 0.484 as you had it To find voids ratio, we can draw the phase diagram of the wet soil / cubic foot air water solid total ************|****************|*************|***************** W 0 11 112 123 lbs V 0.149 11/62.4 112/(2.66*62.4) 1.0 c.f. =0.176 = 0.674 Vv = volume of voids (per cubic ft) = 0.149+0.176 = 0.325 Voids ratio = Vv/Vs = 0.326/0.674 = 0.484 (d) porosity (almost) the same as you had it. porosity n = Vv/Vt = .326 / 1 =0.325

OpenStudy (anonymous):

@mathmate thanks for checking it. Since you have helped me out so much I have become better at these types of problems as you can see. Thanks again for being a great help.

OpenStudy (mathmate):

@ash90 Like everything else, a little practice helps a lot, and that works for me too! :)